# Ravenea: A Benchmark for Multimodal Retrieval-Augmented Visual Culture Understanding

### v1

@@ Appendix I Details of Human Evaluation in the cIC Task
Open Weights Closed Weights Method Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $CCUB_{Rouge-L}$ W/O RAG 18.3 18.7 18.2 21.8 18.2 16.5 17.8 19.6 19.8 16.7 15.9 20.6 15.4 21.4 19.6 Frozen Models SigLIP2-SO/14@384px 18.0 20.0 19.1 23.8 17.2 14.9 19.2 17.4 17.5 15.1 14.9 23.1 14.0 23.6 17.0 CLIP-L/14@224px 18.1 20.9 18.5 23.0 16.8 14.5 18.3 17.2 17.5 15.1 15.8 22.4 14.4 24.7 16.1 Finetuned Models VisualBERT 18.6 19.5 18.3 22.5 16.8 14.8 17.4 17.4 16.9 15.0 16.0 22.5 14.8 24.3 19.0 VL-T5 17.6 19.0 18.1 23.1 17.9 14.4 18.2 17.2 16.9 14.7 15.4 20.8 13.8 23.5 17.4 LLaVA-OneVision-7B 17.3 18.4 17.7 23.0 17.2 14.5 17.6 16.4 17.1 14.6 15.0 20.9 13.1 22.8 17.3 SigLIP2-SO/14@384px 18.4 19.3 19.2 24.5 18.5 14.8 18.4 18.2 17.2 15.2 16.4 22.2 14.3 23.5 18.9 CLIP-L/14@224px 18.1 19.6 18.6 23.9 17.1 14.7 18.0 17.3 17.7 16.0 16.3 21.5 13.7 24.2 17.0 $CCUB_{CIDER}$ W/O RAG 28.1 30.0 27.2 49.2 22.9 16.9 10.5 18.6 35.8 11.2 10.1 49.1 16.7 47.4 34.4 Frozen Models SigLIP2-SO/14@384px 25.2 39.2 20.5 58.5 16.9 7.6 12.4 7.1 6.3 4.7 6.1 62.0 10.1 51.4 17.2 CLIP-L/14@224px 23.8 40.9 18.8 59.6 13.0 6.2 9.6 7.8 4.8 4.7 6.6 56.3 11.1 49.7 14.6 Finetuned Models VisualBERT 26.1 40.2 24.0 60.9 16.3 8.2 10.0 7.5 5.1 6.4 10.2 51.1 15.1 49.5 24.4 VL-T5 24.5 37.0 20.6 51.1 16.0 7.3 10.2 5.0 4.0 6.8 7.5 52.5 10.3 48.4 21.6 LLaVA-OneVision-7B 21.7 38.8 18.5 62.6 13.2 4.8 8.4 5.7 3.4 5.7 4.5 48.9 11.4 42.2 17.6 SigLIP2-SO/14@384px 28.2 40.2 22.8 64.7 17.8 8.4 9.7 5.8 4.4 4.7 7.7 62.3 13.2 54.8 26.5 CLIP-L/14@224px 24.6 35.4 22.1 61.5 15.7 8.9 9.9 6.5 5.8 5.8 7.8 53.1 11.2 51.7 18.3 $CCUB_{BERTScore}$ W/O RAG 54.7 55.4 55.6 55.6 55.7 54.2 52.5 54.5 55.4 54.1 54.8 53.5 52.7 56.6 56.8 Frozen Models SigLIP2-SO/14@384px 54.9 57.2 56.4 57.9 55.2 53.0 55.1 54.8 55.0 52.4 54.0 56.9 52.6 58.7 54.8 CLIP-L/14@224px 54.8 57.6 56.2 57.7 54.8 52.4 54.8 55.1 55.3 52.6 53.7 55.9 52.5 59.4 54.6 Finetuned Models VisualBERT 54.8 57.1 56.5 57.6 54.9 52.8 54.6 54.9 54.4 52.1 53.8 55.7 52.8 58.5 56.2 VL-T5 54.2 56.7 56.0 58.1 54.9 52.6 53.9 54.0 54.4 52.0 53.4 54.4 52.1 57.8 55.7 LLaVA-OneVision-7B 54.2 56.6 55.5 58.2 54.5 52.1 53.9 54.2 54.6 51.6 53.3 55.0 51.8 58.3 54.9 SigLIP2-SO/14@384px 55.2 57.4 57.1 57.9 55.7 53.2 54.7 55.2 54.9 52.5 54.2 56.8 52.6 58.7 56.7 CLIP-L/14@224px 55.0 57.0 56.5 58.2 54.8 53.2 54.6 55.3 55.3 52.9 54.2 56.0 52.3 59.3 55.4 $CCUB_{CLIPScore}$ W/O RAG 19.1 19.1 18.8 19.6 18.3 18.2 18.4 18.7 19.2 18.4 18.4 20.2 18.5 20.2 18.7 Frozen Models SigLIP2-SO/14@384px 19.1 19.3 19.0 19.9 18.9 18.8 19.1 18.8 18.7 18.6 18.5 20.5 18.8 20.0 18.6 CLIP-L/14@224px 19.0 19.2 19.0 20.0 19.0 18.5 19.2 18.8 18.6 18.8 18.4 20.4 18.5 19.6 18.3 Finetuned Models VisualBERT 19.1 19.1 18.9 20.2 18.9 18.9 19.1 18.9 18.8 18.5 18.6 20.3 18.7 19.9 18.5 VL-T5 19.1 19.4 19.1 19.9 19.0 18.8 19.0 19.1 18.7 18.5 18.7 20.4 18.7 19.9 18.4 LLaVA-OneVision-7B 19.1 19.3 18.9 20.1 18.9 18.7 19.3 18.9 18.6 18.5 18.6 20.4 18.8 19.7 18.6 SigLIP2-SO/14@384px 19.1 19.2 18.9 20.1 18.8 18.6 19.1 19.0 18.5 18.6 18.5 20.4 18.8 19.8 18.5 CLIP-L/14@224px 19.0 19.2 18.9 20.0 19.1 18.6 19.2 18.9 18.7 18.6 18.4 20.1 18.8 19.7 18.6
Figure 11: Average improvement across 14 VLMs in different countries with 4 retrievers for cVQA task.
As shown in Figures 12 and 11, VLMs demonstrate varying degrees of performance shifts across cultural contexts in both the cIC and cVQA tasks. In the cIC setting, most fine-tuned retrievers yield noticeable improvements over their original counterparts, indicating the effectiveness of retrieval adaptation. Performance on the cVQA task reveals more nuanced outcomes. While fine-tuned retrievers generally exhibit large performance improvement compared to non-RAG baselines, certain countries, such as Spain and Indonesia, experience exhibit diminishing returns. These discrepancies may stem from the limited presence of culturally representative visual content in the training data. Although targeted image augmentation strategies were employed to alleviate this imbalance, the results suggest that data distribution remains a significant bottleneck. Understanding and addressing such cross-cultural performance disparities in multimodal tasks like cVQA will be an important direction for future work.
Figure 12: Average improvement across 14 VLMs in different countries with 4 retrievers in cIC task.
Appendix K Human Annotation & Evaluation Interface
The annotation interface are shown in Figure 14 and 15. The interface for human evaluation in cIC task is shown in Figure 13.
Figure 13: Human evaluation interface for cIC task.
Figure 14: Human annotation instructions.
(a) Annotation interface (1).
(b) Annotation interface (2).
Figure 15: Human annotation interfaces.
Appendix L Correlation between Automatic Metrics and Human Judgments
To assess the correlation between human preferences and automatic evaluation metrics, we compute Kendall’s Tau rank correlation. Human annotations are segmented according to the output chunks produced by corresponding 14 VLMs with each retriever variant (CaCLIP-based, CLIP-based, and non-RAG). Within each segment, we calculate the selection winning ratio for each retrieval method, yielding a human preference vector formed by concatenating these ratios across all evaluation instances. For the automatic evaluation, we extract BERTScore, CIDEr, ROUGE-L, and CLIPScore for each corresponding retrieval variant and VLM. These scores are similarly concatenated into a metric-based vector. Finally, we compute Kendall’s Tau between the human and metric vectors to quantify the consistency between automatic rankings and human judgments.
Appendix M Prompts
@@
The prompts that we used to collect the documents and two downstream tasks were designed to ensure consistency across different models. One prompt example for cultural caption generation is shown in Table 10. The prompts for both with and without retrieval augmentation for VLMs are shown in Table 12 and 11.
@@
Prompt for data collection GPT-4o SYSTEM: Default
USER:
Generate a culture related caption given the image with around 2 sentences. Please include the name of the thing shown in the picture within the caption if it’s strongly related to the culture. Name the three most culturally relevant entities in the image and attach the name at the end. Follow the format of: caption. entity name1; entity name2; entity name3.
@@
Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
@@
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content and culturally relevant elements of the image. Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph..."). Do not generate multiple options.
@@
Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
The scope of the question is strictly limited to the given image. However, please analyze and incorporate information from both the image and the following document to answer the question.
@@
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content and culturally relevant elements of the image. Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph..."). Do not generate multiple options. Please consider the following context:

### v2

@@ I.3 Vision-language models
Method Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OV-7B $CCUB_{Rouge-L}$ W/O RAG 18.0 19.4 18.1 24.3 18.3 17.1 17.4 18.6 17.0 15.3 14.4 15.2 15.1 23.6 Frozen Models SigLIP2-SO/14@384px 14.9 11.6 16.9 18.7 16.1 14.8 14.0 15.3 16.2 13.0 12.6 12.4 15.0 17.3 CLIP-L/14@224px 15.3 13.6 18.1 19.8 17.0 14.8 13.1 15.1 15.5 12.9 13.8 12.9 15.0 17.3 Finetuned Models VisualBERT 14.6 11.1 16.1 17.4 15.7 14.7 12.2 15.3 15.6 12.9 13.7 11.8 15.8 18.0 VL-T5 14.4 10.6 15.9 18.0 16.8 14.3 13.0 15.1 15.6 12.6 13.1 10.1 14.8 17.2 LLaVA-OV-7B 14.0 10.0 15.5 17.9 16.1 14.4 12.4 14.3 15.8 12.5 12.7 10.2 14.1 16.5 Ravenea-SigLIP 14.8 11.7 16.4 17.3 15.5 15.5 13.3 16.1 15.6 13.2 13.0 13.0 15.1 16.9 Ravenea-CLIP 15.3 12.1 16.7 19.5 18.1 15.6 12.8 15.6 16.1 12.5 13.3 13.6 14.7 18.3 $CCUB_{CIDER}$ W/O RAG 22.2 37.4 26.1 60.7 19.1 11.0 5.7 20.2 10.8 7.5 6.5 6.3 18.0 59.2 Frozen Models SigLIP2-SO/14@384px 8.5 18.3 12.5 17.5 9.8 3.5 2.0 5.6 3.0 1.4 2.0 0.6 13.1 20.5 CLIP-L/14@224px 9.0 21.5 17.0 18.9 11.5 4.3 0.6 3.2 3.3 3.4 2.7 1.6 10.6 18.0 Finetuned Models VisualBERT 9.5 19.3 16.0 19.9 9.2 4.1 0.5 6.0 1.8 3.1 6.1 1.3 18.1 18.6 VL-T5 7.2 16.1 12.6 10.1 8.9 3.2 0.2 3.5 0.7 3.5 3.4 0.9 13.3 17.5 LLaVA-OV-7B 7.1 17.9 10.5 21.6 6.1 0.7 1.2 4.2 0.1 2.4 0.4 1.2 14.4 11.3 Ravenea-SigLIP 8.0 16.8 12.9 12.5 10.0 5.4 0.4 5.3 3.6 1.8 1.8 0.7 13.3 19.0 Ravenea-CLIP 8.5 16.8 11.8 17.3 10.7 5.0 0.1 4.5 3.5 1.3 1.9 1.1 12.6 23.3 $CCUB_{BERTScore}$ W/O RAG 53.6 53.3 55.3 56.2 54.5 53.2 51.7 54.2 53.6 50.1 51.4 51.9 53.5 57.5 Frozen Models SigLIP2-SO/14@384px 50.9 46.7 52.3 54.2 53.1 50.4 48.9 51.8 51.9 48.1 50.3 49.3 51.2 53.7 CLIP-L/14@224px 51.9 49.7 53.1 54.5 53.7 51.6 49.4 52.8 52.5 49.2 50.9 49.8 52.7 55.0 Finetuned Models VisualBERT 50.6 46.6 52.4 53.9 52.8 50.2 48.4 51.9 51.3 47.8 50.1 48.1 51.4 53.5 VL-T5 50.2 46.2 51.9 54.4 52.8 50.0 47.7 51.0 51.3 47.7 49.7 46.8 50.7 52.8 LLaVA-OV-7B 50.2 46.1 51.4 54.5 52.4 49.5 47.7 51.2 51.5 47.3 49.6 47.4 50.4 53.3 Ravenea-SigLIP 51.2 46.7 53.0 52.7 52.8 51.7 49.1 53.0 52.6 48.9 50.5 49.3 51.6 54.1 Ravenea-CLIP 52.5 49.2 54.2 54.9 54.2 52.9 49.3 53.2 53.0 49.0 51.4 51.3 53.4 56.4 $CCUB_{CLIPScore}$ W/O RAG 33.4 33.3 34.1 32.5 33.8 33.4 34.1 34.3 34.5 33.9 33.8 32.4 32.6 32.2 Frozen Models SigLIP2-SO/14@384px 32.7 29.6 32.9 33.4 32.7 33.0 33.8 33.0 33.7 32.1 33.7 31.8 31.9 33.3 CLIP-L/14@224px 32.7 31.4 33.4 33.7 33.2 33.7 35.1 34.1 34.1 33.2 34.1 32.7 32.5 33.6 Finetuned Models VisualBERT 32.7 29.4 32.8 33.7 32.7 33.1 33.8 33.1 33.8 32.0 33.8 31.6 31.8 33.2 VL-T5 32.8 29.7 33.0 33.4 32.8 33.0 33.7 33.3 33.7 32.0 33.9 31.7 31.8 33.2 LLaVA-OV-7B 32.7 29.6 32.8 33.6 32.7 32.9 34.0 33.1 33.6 32.0 33.8 31.7 31.9 33.0 Ravenea-SigLIP 32.7 29.8 33.0 32.1 33.1 33.5 33.8 33.1 33.8 32.7 33.7 31.7 32.0 33.1 Ravenea-CLIP 33.5 31.6 33.4 34.0 33.6 33.7 34.7 34.2 34.2 33.0 33.9 32.8 32.8 33.3
Appendix K Prompts
@@
The prompts that we used to collect the documents and two downstream tasks were designed to ensure consistency across different models. One prompt example for cultural caption generation is shown in Table 13. The prompts for both with and without retrieval augmentation for VLMs are shown in Table 15 and 14.
@@
Prompt for data collection GPT-4o SYSTEM: Default USER:[Uncaptioned image] Generate a culture related caption given the image with around 2 sentences. Please include the name of the thing shown in the picture within the caption if it’s strongly related to the culture. Name the three most culturally relevant entities in the image and attach the name at the end. Follow the format of: caption. entity name1; entity name2; entity name3.
@@
Answer the following multiple choice question for the image. The last line must be of the following format: ’Answer: \$LETTER’ (without quotes) where LETTER must be one of A, B, C, or D.
@@
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content, culturally-relevant elements, and the country’s name. Avoid phrases like ’This image’ or ’The image’.
@@
Answer the following multiple choice question for the image. The last line must be of the following format: ’Answer: \$LETTER’ (without quotes) where LETTER must be one of A, B, C, or D.
Use the shared culture between the image and the following document to answer the question.
@@
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content, culturally-relevant elements, and the country’s name. Avoid phrases like ’This image’ or ’The image’. Please consider the following DOCUMENT as supplementary material for the image. Mention the country’s name of the culture.
Document:
@@
Appendix L Human Annotation & Evaluation Interface
The annotation interfaces are shown in Figure 13 and 14. The interface for human evaluation in cIC task is shown in Figure 12.
Figure 12: Human evaluation interface for cIC task.
Figure 13: Human annotation instructions.
(a) Annotation interface (1).
(b) Annotation interface (2).
Figure 14: Human annotation interfaces.